# Halilsoy and Chandrasekhar standing gravitational waves in the linear approximation

### v1

@@ V Summary
We discovered that the Halilsoy (1) solution and the Chandrasekhar solution (3) are, in the linear approximation, gravitational analogues of of two types of electromagnetic polarization standing waves [10]. This analogy holds strictly far from the symmetry center, where the fluctuations in C-energy vanish. Furthermore, we have clarified in which sense the solutions under investigation are standing gravitational waves near the symmetry center — the circumference of the test tube’s cross-section exhibits characteristic for standing waves behaviour, with fixed positions of nodes and antinodes.
In the light of our results, the less restrictive definition of standing gravitational waves provided by Stephani [1] — according to which C-energy must be constant only on average in time — appears to be sufficient. Although the Halilsoy solution does not satisfy the more restrictive Chandrasekhar definition [2], it still exhibits typical standing waves behaviour in the weak-field approximation.

### v2

@@ V Summary
We discovered that the Halilsoy (1) solution and the Chandrasekhar solution (3) are, in the linear approximation, gravitational analogs of of two types of electromagnetic polarization standing waves [10]. This analogy holds strictly far from the symmetry center, where the fluctuations in C-energy vanish. Furthermore, we have clarified in which sense the solutions under investigation are standing gravitational waves near the symmetry center—the circumference of the test tube’s cross section exhibits characteristic for standing waves behavior, with fixed positions of nodes and antinodes.
In light of our results, the less restrictive definition of standing gravitational waves provided by Stephani [1]—according to which C-energy must be constant only on average in time—appears to be sufficient. Although the Halilsoy solution does not satisfy the more restrictive Chandrasekhar definition [2], it still exhibits typical standing waves behavior in the weak-field approximation.
@@ Acknowledgments
Data availability
No data were created or analyzed in this study.